Wednesday, 19 August

Wednesday, 19 August

Short Communications and Poster Sessions

Short Communications: Wednesday, 19 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 02

Rings, Chair: Toma Albu

MA 042 Hyoung J. Ko Some complete intersection and their minimal prime ideals (p.20)

Takasi Nagahara

Classification of rings satisfying some generalized Herstein’s condition (p.24)

Sylvia Chin-Pi Lu

The Zariski topology on the prime spectrum of a module (p.22)

Sect.: 03

Computational Number Theory, Chair: Michael E. Pohst

MA 311 Claus Fieker On Constructive Class Field Theory (p.36)

István Gaál

Power integral bases in algebraic number fields (p.36)

Jürgen Klüners

Explicit Galois realization of transitive groups of degree up to 15 (p.40)

Sect.: 08

Function Spaces and Operators, Chair: Jose Bonet

MA 041 Silke Holtmanns An operator representation for weighted spaces of vector valued holomorphic functions (p.130)

Klaus D. Bierstedt

An operator representation for weighted inductive limits of spaces of vector valued holomorphic functions (p.117)

Jose Bonet

Hypercyclic operators on non-normable Fréchet spaces (p.118)

Sect.: 08

Iteration, Composition, Clifford Structure, Chair: Mario Bonk

MA 309 Chung-Cun Yang Some recent results on factorization of meromorphic functions (p.165)

Eugen J. Ionascu

Wandering sets for a class of Borel isomorphisms of [0, 1) (joint work with Dr. Edward Azoff, UGA, Athens, USA) (p.131)

Enrique Ramirez-De-Arellano

Almost-Clifford structures on manifolds (p.150)

Sect.: 08

Fractional Differentiation, Chair: A. Eduardo Gatto

MA 043 Evelyn Buckwar Application of the Lie group of scaling transformations to the fractional diffusionwave equation (p.119)

Yuri Luchko

Lie group analysis of a partial differential equation of fractional order (p.140)

A. Eduardo Gatto

Fractional Calculus, Sobolev Spaces, and Related Rules on Spaces of Homogeneous Type (p.124)

Sect.: 10

General Theory, Chair: Massimo Lanza de Cristoforis

H 111 Wenxiong Chen High Order Conformally Invariant Equations (p.200)

Paweł Domański

Splitting of differential complexes (p.201)

Massimo Lanza de Cristoforis

Differentiability Properties of the Autonomous Composition Operator in Sobolev Spaces (p.213)

Sect.: 10

Elliptic Equations I, Chair: Florin C. Muscalu

H 110 Evgeni A. Kalita New class of a priori estimates for nonlinear elliptic equations and systems of high order

(p.208)

Vladimir A. Kondratiev

On Properties of Positive Solutions of Elliptic Equations (p.210)

Florin C. Muscalu

A Harnack Type Theorem for Elliptic Systems with L^{∞} Coefficients (p.214)

Sect.: 12

Branching Processes, Chair: Hermann Dinges

H 107 Anton Wakolbinger Critical branching random walks in a random medium (p.272)

Aubrey Wulfsohn

Branching diffusions with random offspring distributions (p.272)

Klaus Fleischmann

Super-Brownian motion with a super-Brownian catalyst (p.250)

Sect.: 12

Data Analysis, Chair: Joerg Polzehl

H 1029 Osvaldo Marrero Statistical analysis of seasonality in multiple groups (p.260)

Fernando Rosado

On the Outlier Problem for Structured Data (p.268)

Vera Pawlowsky-Glahn

Distances for multivariate spatial data analysis (p.265)

Sect.: 15

Ordinary Differential Equations I, Chair: Alexander Ostermann

H 2013 Gennady Yu. Kulikov How to control the global error of Runge-Kutta methods (p.305)

Alexander Ostermann

The unstable manifold of an equilibrium point under discretization (p.308)

George Micula

Polynomial spline functions of even degree approximating the solution of differential equations

Sect.: 16

Phase Transition and Free Boundary Value Problems, Chair: Jürgen Socolowsky

H 106 Nikolaus Bubner Phase Transitions in Shape Memory Alloys

(p.320)

Olaf Klein

Approximation results for phase–field systems (p.326)

Jürgen Socolowsky

On a perturbed two-layer free surface flow down an inclined plane (p.332)

Sect.: 16

Coding and Image Compression, Chair: John K. Karlof

H 112 Alexander P. Petukhov Recursive wavelets and image compression

(p.330)

John K. Karlof

A Coding Scheme for the Gaussian Multiple Access Channel (p.325)

Sect.: 17

Applied Optimal Control, Chair: Fredi Tröltzsch

H 2033 Hans Josef Pesch Optimal Control of Large Scale DAE Systems in Chemical Engineering (p.347)

Fabian Flores-Bazan

Optimal solutions to some variational problems arising in Economics (p.341)

Bernd Kugelmann

Robust and Optimal Feedback-Control in Flight Dynamics (p.344)

Short Communications: Wednesday, 19 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 02

Algebras I, Chair: Alexandre V. Mikhalev

MA 042 Askar Dzhumadil’daev Cohomologies of right-symmetric algebras and Novikov algebras (p.17)

Ekaterina A. Vasilieva

Involutive Bases of Free Algebras (p.31)

Alexander A Totok

Actions of Hopf algebras on noncommutative algebras

Sect.: 03

Additive Number Theory, Chair: Doytchin Tolev

MA 311 Mariano Garcia New type (i, 1) amicable pairs with greatest common factor a power of 2 or a prime times a power of 2 (p.37)

Nazir M. Khan

The Numbers that Can Be Represented as the Sum of Four Squares (One Repeating Itself Three Times), in Three Different Ways (p.39)

Doytchin Tolev

Additive problem with prime numbers of a special type (p.46)

Sect.: 08

C*-Algebras and Modules, Chair: Alexander Ya. Helemskii

MA 041 Michael J. Gruber Non-commutative Bloch theory (p.128)

Evgenij V. Troitsky

Kuiper and Dixmier-Douady theorems for C^{∗}-Hilbert modules and K-theory (p.160)

Alexander Ya. Helemskii

Description of projective Hilbert modules over C^{∗}-algebras (p.129)

Sect.: 08

Spaces of Analytic Functions, Chair: Peter Duren

MA 309 Mitsuo Morimoto Reproducing kernels related to the complex sphere (p.142)

Romeo Meštrović

Topologies on Some F -algebras of Holomorphic Functions (p.141)

William E. Hornor

Properties of isometry-inducing maps of the unit disc (p.130)

Sect.: 08

Wavelet Transforms, Chair: Boris Rubin

MA 043 Charles N. Moore Distribution function inequalities for sums of wavelets (p.142)

John Schmeelk

Wavelet Transforms in Generalized Fock Spaces (p.154)

Boris Rubin

Fractional Calculus and Wavelet Transforms in Integral Geometry (p.152)

Sect.: 10

A Priori Estimates & Asymptotics, Chair: Shuji Watanabe

H 111 Sergey Frolovitchev Multiplicative solution asymptotics for Kolmogorov-Feller equation and asymptotics of the boundary value problems

(p.203)

Annegret Glitzky

Electro–reaction–diffusion systems in heterostructures (p.204)

Shuji Watanabe

An embedding theorem of Sobolev type related to a generalized Fourier transform and applications to partial differential equations with singularity (p.224)

Sect.: 10

Applied Fluids, Chair: Lizabeth V. Rachele

H 110 Olaf Schmidtmann Stability Problems in Turbo Compressors

(p.220)

Lizabeth V. Rachele

Inverse Problems for Elastic Media (p.218)

Sect.: 12

Random Walks, Discrete Structures, Chair: Klaus Fleischmann

H 107 Boris L. Granovsky Spectral gap of certain continuous-time Markov processes (p.252)

Rainer Siegmund-Schultze

Level Crossing Probabilities for Random Walks (p.269)

Sect.: 12

Distribution Theory, Chair: Lutz Mattner

H 1029 Tea-Yuan Hwang On some characterizations of population distributions (p.254)

Papa Ngom

Aspect bayésien des a-distributions et application (p.263)

Lutz Mattner

What are cumulants? (p.261)

Sect.: 15

Ordinary Differential Equations II, Chair: Tatiana Anatolyevna Vasilyeva

H 2013 Felix Schwitzer Linearly implicit Runge-Kutta discretizations of semilinear parabolic problems near equilibria (p.312)

Sergei Yu. Pilyugin

Shadowing and perturbation of attractors in discretized parabolic equations (p.310)

Tatiana Anatolyevna Vasilyeva

Twice implicit scheme with accuracy control and it application to chemical kinetics problems (p.316)

Sect.: 16

Numerics for Wave Propagation, Chair: Georg Hebermehl

H 106 Rainer Schlundt On the Computation of Systems of Linear Algebraic Equations for Monolithic Microwave Integrated Circuits (p.332)

Hüsnü Ata Erbay

A numerical study of the complex modified Korteweg-de Vries equation (p.323)

Georg Hebermehl

On the Computation of Eigen Modes for Lossy Monolithic Microwave Integrated Circuits (p.324)

Sect.: 17

Distributed Parameter Systems I, Chair: Dan Tiba

H 2033 Pierre, F. Aubert Shape optimisation of geometrically nonlinear 3d beams and shells (p.336)

Raja Dziri

Dynamical shape optimazation in Navier-Stokes equations (p.339)

Michael Hinze

On the construction of feedback controllers for parabolic equations with an application to the control of fluid flow

(p.343)

Short Communications: Wednesday, 19 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 02

Semigroups, Chair: Hans-Jürgen Vogel

MA 311 Olga Macedonska On semigroup identities in groups (p.22)

Boris M. Schein

A new construction for free inverse semigroups (p.28)

Vučić Dašić

Hyperrings and hypernear-rings of the multivalied mappings (p.16)

Sect.: 02

Topological Algebra, Chair: Karl-Heinrich Hofmann

MA 042 Nouzha El Yacoubi J-Generalized topological divisors of zero in topological Jordan algebras (p.18)

Rumara Jewett

Almost Convexity of a Certain Nilpotent Group (p.19)

Mary Nailevna Sabitova

The automorphisms of affine spaces and polynomial manifolds (p.27)

Sect.: 05

Kähler Manifolds, Chair: Bill Watson

MA 041 Keizo Hasegawa A class of compact Kählerian solvmanifolds and a general conjecture (p.71)

Toyoko Kashiwada

The integrability of almost generalized Hopf manifolds (p.72)

Bill Watson

The Goldberg Conjecture is true for the four-dimensional total space of an almost Hermitian submersion (p.84)

Sect.: 06

Group Actions, Chair: Reinhard E. Schultz

H 112 Charles B. Thomas Contact structure and spherical space forms (p.98)

Harald Biller

Actions of large compact Lie groups on spheres (p.85)

Takashi Tsuboi

On the group of area preserving homeomorphisms (p.99)

Sect.: 09

Symmetry and Bifurcation I, Chair: Karin Gatermann

H 111 Oluwole Daniel Makinde Numerical Study of Bifurcation for Steady Flow in a Diverging Tube (p.182)

Dmitry Turaev

On infinitely degenerate dynamical systems (p.194)

Karin Gatermann

Orbit space reduction (p.175)

Sect.: 09

Dynamical Systems on Surfaces, Chair: Raphael Albrecht

H 110 Victoria Rayskin Degenerate homoclinic crossings (p.187)

Shobh N. Singh

Solutions of boundary value problems by double Laplace transformations (p.192)

Iuri M. Zaitsev

Lyapunov-type vector-function for nonlinear dynamic system (p.196)

Sect.: 13

Geometry I, Chair: Jörg Rambau

H 106 Dietmar Cieslik The Steiner Ratio of Metric Spaces

(p.275)

Michael Joswig

On Elation Laguerre Planes (p.277)

Alexandre A. Makhnev

Partial geometries and the Krein condition (p.282)

Sect.: 15

Wavelets I (organized by Andreas Rathsfeld), Chair: Wayne M. Lawton

H 2033 Jacques Liandrat Wavelets, from theory to efficients numerical algorithms (p.306)

Andreas Rieder

Embedding and a-priori adaptivity (p.311)

Pierre Gilles Lemarié-Rieusset

Some remarks on divergence free vector wavelets (p.306)

Sect.: 15

Nonlinear Operator Equations, Chair: Jose A. Ezquerro

H 2013 Jose A. Ezquerro A semilocal convergence result for Secanttype methods under Hölder continuity conditions (p.299)

Anatolii Borisovich Bakushinskii

Iterative methods for the solution of nonlinear operator equations without property of the regularity (p.296)

Victor D. Didenko

On Peculiarities of Approximation Methods for Singular Integral Equations with Conjugation (p.298)

Short Communications: Wednesday, 19 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 02

Groups I, Chair: Wujie Shi

MA 311 Herbert Abels Reductive groups as metric spaces (p.12)

Oleg I. Tavgen

The free product of cyclic groups in the group PSL_{2}(ℂ) (p.30)

Matthias Trittler

Normalteiler symplektischer Gruppen (p.30)

Sect.: 02

Semirings/groups and Algebras, Chair: Boris M. Schein

MA 042 Irena I. Artamonova We consider semirings with 1 ≠ 0 (p.12)

Godwin Ekhaguere

Algebraic representation theory of partial algebras (p.17)

Miroslav, D. Ćirić

Atomic decompositions of algebras and structures

Sect.: 05

Topology of Manifolds, Chair: Sebastian Goette

MA 041 Christina Sormani The Topology of Noncompact Manifolds with Positive Ricci Curvature (p.80)

Gerald M. Lodder

Leibniz Cohomology for Differentiable Manifolds (p.75)

Sebastian Goette

Equivariant η-invariants and η-forms (p.70)

Sect.: 05

Miscellaneous Topics, Chair: Zheng-Chao Han

MA 043 Margaret F. Symington Symplectic rational blowdowns (p.81)

Zheng-Chao Han

New existence results on conformal metrics with constant interior scalar curvature and constant boundary mean curvature (p.71)

Sect.: 06

Knots, Chair: Hitoshi Murakami

H 112 Alan H. Durfee Polynomial representations of knots (p.87)

Hugh R. Morton

The Burau matrix and Fiedler’s invariant for closed braids (p.94)

Vladimir V. Vershinin

Generalizations of Braids: Homology and Vassiliev Invariants (p.100)

Sect.: 09

Symmetry and Bifurcation II, Chair: Reiner Lauterbach

H 111 Patrick Bonckaert Convergence problems of normalizing transformations with symmetry (p.172)

Claudia Wulff

Spiral Waves and Euclidean Symmetry (p.194)

Reiner Lauterbach

Intermittent dynamics and symmetry (p.180)

Sect.: 09

Invariant Measures and Random Dynamics I, Chair: Gerhard Keller

H 110 Alexander Loskutov Dynamics of one-dimensional nonhomogeneous coupled map lattices (p.181)

Roberto Markarian

Invariant measures for Anosov maps with small holes (p.182)

Nikolas Efstathiou Sofronidis

The Topological Entropy is continuous on a second category set in Diff^{r}(M), whenever M is a compact manifold and r is a non-negative integer (p.193)

Sect.: 11

Many Particle Systems, Chair: Gennady Shchepanuk

H 107 Valeri V. Kucherenko Normal forms of quadratic boson operators and theirs spectrum (p.238)

Christoph Bandt

The geometry of a parameter space for interacting particle systems (p.229)

Gennady Shchepanuk

Almost L_{2}-exponential mixing for classical continuous systems (p.245)

Sect.: 13

Geometry II, Chair: Günter M. Ziegler

H 106 Min Yan Stratified Approach in Combinatorics

(p.287)

Jörg Rambau

A new View on the Cayley Trick (p.283)

Masahiro Hachimori

Non-constructible triangulations of balls (p.277)

Sect.: 15

Interpolation, Chair: Raquel R. Martinez

H 2033 Raquel R. Martinez Composite Methods for Interpolating Functions (p.307)

Wolfgang Volk

Quasi-interpolation with Arbitrarily Often Differentiable Functions (p.316)

Yurij Kaluch I.

Mathematical Problems of Condition Prognosing of Land Slopes and Slides

(p.303)

Poster Session: Wednesday, 19 August, 14:00–15:45

Sect.: 04

Algebraic Geometry

MA 144 Jacqueline Arancibia Canonical Curves in P^{3} (p.49)

Isidro P. Nieto

H_{2,2}-Kummer surfaces associated to polarized abelian surfaces of type (1,3)

(p.58)

Niranjan Ramachandran

A conjecture of Deligne (p.59)

Rudi Weikard Algebro-geometric Differential Operators

(p.62)

Katsuaki Yoshida

Klein’s surface of genus three and associated theta constants (p.62)

Rupert W.T. Yu

Lattice polytopes for certain Demazure modules of sl_{n+1} (p.63)

Sect.: 08  Analysis

MA 144 Aldric L. Brown Operator convex functions of two variables

(p.118)

Mao-Ting Chien

A parametrization of the c-numerical ranges of certain matrices (p.120)

William C. Connett

Measure algebras that have oblate spheroidal wave functions as characters

(p.121)

Sonja Mouton Some spectral properties in ordered Banach algebras (p.143)

Marjatta Näätänen

Arithmetic Fuchsian groups of signature (0; a, b, c, oo) (p.143)

Samyon R. Nasyrov

Robin capacity and lift of infinitely thin airfoils (p.144)

David L. Salinger Lifting the Difference Spectrum (p.153)

Ajit Iqbal Singh

Maximal quotient rings of algebras of functions and operators (p.156)

Arne Stray

Polynomial approximation of analytic functions with finite Dirichlet integral (p.158)

Chia-Chi Tung A Note on the Hilbert Nullstellen Satz

(p.161)

Dmitri Zaitsev

On holomorphic boundary extensions, web theory, and invariants of Grassmanians

(p.165)

Jaroslav Zemánek

Powers and resolvents (p.167)

Poster Session: Wednesday, 19 August, 16:15–18:00

Sect.: 16

Applications

MA 144 Attia A. Ashour Solution for Certain Three Dimensional Boundary Value Problems with Mixed Boundaries (p.320)

Wolfgang Dreyer

Evolution of Shock Structures in Gases and Solids (p.322)

Rolf Hünlich

Modelling and simulation in the technology of power devices (p.324)

Sophia R.-J. Jang The dynamics of a variable-yield nutrientplankton model (p.325)

Hans-Christoph Kaiser

Modeling of QW–Lasers Including Schrödinger–Poisson Systems

(p.325)

Carlos C. Moreira

Self-organized model for the genetic code (p.330)

Alexandra E. Rodkina On stochastic functional-differential equations and its applications (p.331)

Elena Shchepanika

Duck-trajectories in combustion problems (p.332)

Krzysztof Wilmański

Dynamic Singularities in Porous and Granular Materials (p.334)

Hajime Yamashita Approximate Analysis of Fuzzy Graph and its Application (p.334)